# Extracting Interpretable Logic Rules from Graph Neural Networks

### v1

@@ 3.3 Grounding $\phi_{M}$ into the Input Space
3.4 Extensions
Inference with LogicXGNN
Unlike a graph neural network $M$, which classifies based on the highest logit, its rule-based equivalent $\phi_{M}$ classifies an instance as belonging to a class if two conditions are satisfied simultaneously: 1) The instance matches that class’s (discriminative) rules, and 2) it does not match any other class’s rules. On the other hand, if an instance matches the rules of multiple classes, the model reports the conditions instead of making a definitive choice. This approach is safer, as $\phi_{M}$ only makes decisions when confident and provides the reasoning—matched rules—for human inspection in ambiguous cases.
LogicXGNN as Generative Model
As LogicXGNN makes each decision-making step in GNNs transparent with descriptive rules, we can leverage LogicXGNN as a generative model for creating graph instances in a controlled manner. First, we sample a conjunctive form from $\bar{\phi}_{M}$, then combine it with the connectivity pattern $\psi_{M}$ to form a graph consisting of predicates, and finally, assign node features based on the grounding rules associated with the corresponding predicates. We provide more details in Appendix A.
Table 1: Results of the cardinality of the extracted rules $\phi_{M}$, along with a comparison of the train and test accuracies fo $\hat{\phi}_{M}$, the original GNN $M$, and GLGExplainer. The columns labeled $|P|$, $|\bar{\phi}_{M}^{0}|$, $|\bar{\phi}_{M}^{1}|$, $|\hat{\phi}_{M}^{0}|$, and $|\hat{\phi}_{M}^{1}|$ report the number of predicates, as well as the number of conjunction clauses in the descriptive rules $\bar{\phi}_{M}$ and discriminative rules $\hat{\phi}_{M}$ for classes 0 and 1, respectively.
The cardinality of $\phi_{M}$ Train accuracy Test accuracy
Dataset $|P|$ $|\bar{\phi}_{M}^{0}|$ $|\bar{\phi}_{M}^{1}|$ $|\hat{\phi}_{M}^{0}|$ $|\hat{\phi}_{M}^{1}|$
$M$
$\hat{\phi}_{M}$
GLG.
$M$
$\hat{\phi}_{M}$
GLG.
HIN 2293 426 425 146 62 89.12 93.83 50.04 86.93 85.23 49.99
BBBP 254 146 1044 20 22 81.18 87.37 45.18 81.13 81.37 42.90
Mutagenicity 314 827 1015 198 167 78.12 84.03 56.78 76.04 76.50 54.26
Figure 3: Selected knowledge discovered by LogicXGNN. The rows display results for the BBBP and MUTAG datasets, respectively.
@@ 4 Evaluation
In this section, we conduct experimental evaluations on real-world datasets to address the following research questions:
@@
    How effective is $\phi_{M}$ as a classification tool? Does it have the potential to replace the original GNN $M$?
@@
Benchmarks and Baselines
We use the following graph classification benchmarks: (1) The Mutagenicity (MUTAG) dataset consists of 4337 molecule graphs, each labeled based on their mutagenic effect on the Gram-negative bacterium Salmonella typhimurium (Debnath et al., 1991). (2) The BBBP dataset consists of 2039 molecule graphs, each labeled based on their ability to penetrate the blood-brain barrier (Wu et al., 2018). (3) The Hospital Interaction Network (HIN) dataset comprises 1760 graphs representing face-to-face interactions between doctors and nurses. The task is to classify whether the ego in each ego network is a doctor or a nurse (Azzolin et al., 2023). For the baseline approaches, we employ a generation-based method, XGNN (Yuan et al., 2020), as well as two concept-based methods, GCNeuron (Xuanyuan et al., 2023) and GLGExplainer (Azzolin et al., 2023). All these methods are considered state-of-the-art.
Setup
All experiments are conducted on an Ubuntu 22.04 LTS machine with 32 GB of RAM and a 2.7 GHz processor. We use a 2-layer GNN with a hidden dimension of 32 for all datasets. To demonstrate the model-agnostic nature of LogicXGNN, we employ graph convolutional networks for BBBP and HIN, and GraphSAGE (Hamilton et al., 2017) for MUTAG. All datasets are split into train-test sets with an 8:2 ratio. Additionally, we use the CART algorithm for the decision trees in LogicXGNN.
Figure 4: Selected examples generated by descriptive rules $\bar{\phi}_{M}$ and explanation graphs generated by XGNN. The rows display results for BBBP, MUTAG, and HIN, respectively. Each cell under LogicXGNN includes the original graphs, their descriptive rules, and a newly generated instance based on these rules.
@@ 4.1 How effective is $\phi_{M}$ as a classification tool?
We report the cardinality of $\phi_{M}$, the rules extracted by LogicXGNN from the original GNN $M$, and compare its classification accuracy on both the train and test sets with $M$ and baseline approaches in Table 1.
Notably, our rule-based model $\hat{\phi}_{M}$ outperforms its neural counterpart $M$ on the BBBP and MUTAG datasets, despite being derived from it. While $\hat{\phi}_{M}$ shows a slight tendency to overfit the training data, it generalizes quite well on the test data, achieving nearly identical performance to the original model. For example, on the BBBP dataset, $\hat{\phi}_{M}$ achieves an accuracy of $81.37\%$ on the test set, marginally outperforming the $81.13\%$ accuracy of the original GNN $M$. This suggests that $\hat{\phi}_{M}$ successfully captures the underlying decision-making logic of $M$, demonstrating its potential to replace neural models in many settings where high fidelity and interpretability are critical. Having a white-box model like $\hat{\phi}_{M}$ offers many significant advantages. For instance, unlike black-box models, it provides reasoning in the form of interpretable logic rules rather than merely producing predictions. More importantly, it enables users to inspect where the model may make errors, allowing for targeted improvements by incorporating human prior knowledge to further enhance the model’s performance. For instance, the number of carbon rings influences mutagenicity, yet this information is not preserved in the global_mean_pooling operation of GNNs. In contrast, our rule-based model can explicitly incorporate this factor with a simple patch. We leave this direction for future work.
As for the baselines, GCneuron lacks discriminative power, as it generates explanations based solely on the input class. Specifically, providing different target classes for the same input yields entirely different explanations. On the other hand, GLGExplainer underestimates the rich combinatorial complexity of graphs and tends to produce overly simplistic rules that fail to capture the variety of data patterns. For instance, GLGExplainer uses only two predicates, $p_{0}$ and $p_{1}$, to represent explanations for classes 0 and 1 in the Mutagenicity dataset. However, to accurately distinguish between these classes, our approach generates disjunctive normal forms (DNF) of sizes 198 and 167 for classes 0 and 1, respectively. As a result, GLGExplainer’s performance is similar to a random classifier.
@@ 4.2 What knowledge can we derive using $\phi_{M}$?
We ground the discriminative rules $\hat{\phi}_{M}$ and present the selective knowledge extracted from $M$ for the BBBP and MUTAG datasets in Figure 3. These findings are accurate from the perspectives of chemistry and molecular science, and they effectively explain the classification results. For instance, we recognize that molecules containing oxygen-rich groups are less likely to cross the Blood-Brain Barrier (class 0). This is due to their increased hydrophilicity, reduced lipophilicity, and higher likelihood of being recognized by efflux transporters (Pardridge, 2005).
Regarding mutagenicity, we find that molecules containing a methyl group (-$\text{-CH}_{3}-$) are non-mutagenic (class 1). This is because the methyl group is generally less reactive and does not readily participate in chemical reactions that could lead to DNA damage or mutagenesis (Jin & Robertson, 2012). Additionally, we find that the combination of an aromatic or heteroaromatic ring with a nitro group is crucial for mutagenic activity (class 0), consistent with the findings of the original paper (Debnath et al., 1991). In contrast, existing approaches like GCNeuron and GLGExplainer treat nitro groups and carbon rings as isolated entities, overlooking the critical synergy between them.
@@ 4.3 How effective is $\phi_{M}$ as a generative model?
We present selected examples generated by descriptive rules $\bar{\phi}_{M}$ and explanation graphs generated by the baseline approach, XGNN, in Figure 5. As the descriptive rules make each decision-making step in GNNs transparent, we can accurately model the underlying data distribution, enabling the generation of new instances in a fully controlled manner. Specifically, we can select certain conjugate clauses from the descriptive rules, particular connection patterns, and preferences in predicate grounding during the generation process. For instance, we select the conjugate clause $p_{41}\land p_{6}\land p_{15}$ with the following connection pattern: $p_{15}$ is connected to $p_{61}$, $p_{61}$ is connected to $p_{6},p_{6},p_{15}$, and $p_{6}$ is connected to $p_{6},p_{61}$. Given the grounding rule that $p_{41},p_{6},$ and $p_{15}$ all encode motifs composed of carbon atoms, we successfully reconstruct the original molecule.
In contrast, XGNN learns a reinforcement learning agent that is inherently black-boxed, generating graphs that do not align with the actual data distribution. For example, it generates bipartite structures that lack molecular relevance, whereas our approach learns the underlying rules and adheres to them during the generation process. As a result, $\bar{\phi}_{M}$ can not only reconstruct original instances but also generate new ones that preserve similar structures and key properties.
We envision adding more flexibility to the generation process, such as randomly combining conjugate clauses from descriptive rules, which could lead to even more interesting results. This approach may enable newly generated molecular instances to inherit key properties from multiple known molecules. In summary, LogicXGNN demonstrates significant potential in fields such as drug design, and we will further explore these possibilities in future work.

### v2

@@ 3.3 Grounding $\phi_{M}$ into the Input Space
3.4 Analysis
Computational complexity
Our approach models message passing at each node in the dataset to identify predicates. In the first step, we extract activation patterns from pretrained GNNs and compute graph hashes on nodes’ local neighborhoods. This step runs independently of the GNN size, with a complexity of $O(|\mathcal{V}|)$. Second, determining the logical structure involves constructing a binary matrix of size (number of predicates) $\times$ (number of graphs), resulting in a complexity of $O(|\mathcal{V}||\mathcal{G}|)$. Finally, grounding predicates—by selecting features or performing decision tree analysis for each predicate—has a complexity of $O(|\mathcal{V}|)$. We report empirical runtime performance in Table 1.
Generalization across different GNN architectures
We show the theoretical generalizability of LogicXGNN to any GNN architecture. First, we model stacked message-passing computations using hidden predicates (activation patterns and local subgraphs)—an architecture-agnostic formulation. Next, we generate logic rules through binary matrix construction and decision tree analysis, maintaining architecture independence. Finally, we ground predicates by linking them to input features via decision trees, requiring no GNN-specific details. Empirical evidence is provided in Appendix A.6.
@@ 4 Evaluation
In this section, we conduct experimental evaluations on real-world benchmark datasets to address the following research questions:
@@
    How effective is $\phi_{M}$ as a classification tool, especially compared to the original GNN $M$?
@@
Baselines
Consistent with prior work [30, 3, 2], we focus our comparison on global explanation methods, excluding local approaches such as GNNExplainer [31] due to their differing scope. To benchmark our approach across different functionalities, we use rule-based methods—GraphTrail [2] and GLGExplainer [3]—as baselines for evaluating classification performance and knowledge extraction. We also include the generation-based method XGNN [32] to benchmark our graph generation ability. All of these methods are considered state-of-the-art. A summary of popular explanation methods and their supported functionalities is provided in Table 2 (Appendix A.1).
Datasets and Experimental Setup
We use a diverse set of graph classification benchmarks commonly employed in GNN explanation research, with detailed descriptions provided in Appendix A.2. Among these, Mutagenicity [6], NCI1 [25], and BBBP [27] are molecular graph datasets representing chemical compounds; BAMultiShapes [31] is a synthetic dataset focused on structured geometric patterns; and IMDB-BINARY [16] is a social network dataset. To demonstrate the model-agnostic nature of LogicXGNN, we evaluate it using GCN [10] for BAMultiShapes and BBBP, GraphSAGE [8] for Mutagenicity, GIN [29] for NCI1, and GAT [23] for IMDB-BINARY. Evaluation results for each dataset across various GNN architectures are reported in Table 11. Further details on the experimental setup, including GNN training and baseline implementations, are provided in Appendix A.3.
Table 1: Classification accuracy (%) and runtime (seconds) of various explanation methods on benchmark datasets. The first row reports the original GNNs’ accuracy (Runtime is omitted for GNNs as comparisons are only relevant for explanation methods). Subsequent rows present results for explanation methods, with the highest accuracy and fastest runtime among explanation methods highlighted in bold. Additional evaluations across three random seeds are provided in Appendix A.4.
BAShapes BBBP Mutagenicity NCI1 IMDB
Method Acc. Time Acc. Time Acc. Time Acc. Time Acc. Time
GNN 80.50 81.62 76.27 76.28 71.00
GLG 57.50 312.7 52.45 359.8 57.95 726.4 53.41 875.74 53.00 329.43
G-Trail 82.00 2,543.2 82.11 5,647.9 65.90 20,049.3 66.42 24,067.5 57.50 1,073.4
$\phi_{M}$ (Ours) 90.00 45.3 83.58 52.8 77.76 103.6 76.03 135.4 65.00 69.5
@@ 4.1 How effective is $\phi_{M}$ as a classification tool?
To answer this question, we report the test set classification accuracy of LogicXGNN ($\phi_{M}$) against baseline approaches and the original GNN $M$ in Table 1.
Notably, LogicXGNN consistently outperforms both baseline approaches—GraphTrail and GLGExplainer—by a significant margin across all benchmarks. This performance gap can be attributed to fundamental differences in design philosophy. GLGExplainer applies clustering algorithms to local explanations to construct global explanations, while GraphTrail uses symbolic regression to fit a surrogate rule model on subgraph-level concepts as its explanation. In contrast, LogicXGNN is designed to model the actual computational flow in GNNs—such as message passing and information pooling—naturally deriving DNF rules as explanations. As a result, both GraphTrail and GLGExplainer tend to underestimate the complexity of the GNNs’ decision-making process, an issue to which LogicXGNN is less susceptible. In terms of runtime performance, LogicXGNN achieves one to two orders of magnitude speedup—approximately 10 to 100 times faster—compared to GLGExplainer and GraphTrail, by leveraging efficient decision tree algorithms and graph traversal. This eliminates the need for training local surrogate models or performing symbolic regression, significantly reducing computational overhead.
Another interesting observation is that $\phi_{M}$ can even outperform the original model on certain benchmarks. For instance, on the BAMultiShapes dataset, $\phi_{M}$ achieves a test accuracy of $90.00\%$, significantly surpassing the $80.50\%$ accuracy of the original GNN model $M$. This superior performance is also observed and validated in more challenging out-of-distribution scenarios, as discussed in Appendix A.7. Upon closer inspection, we find that LogicXGNN generally excels on datasets with well-structured, domain-specific patterns, as is often the case with molecular datasets. We believe such datasets may inherently align with a logic-rule-driven structure, allowing GNNs trained on them to exhibit decision-making patterns that can be effectively captured by LogicXGNN. This demonstrates LogicXGNN’s potential to replace neural models in scenarios where both high fidelity and interpretability are critical.
However, for graphs with more complex node connections and weaker structural regularities, such as IMDB-BINARY, a performance gap remains. We aim to address this limitation in future work.
@@ 4.2 What knowledge can we derive using $\phi_{M}$?
Figure 2: Chemical knowledge extracted by different explanation methods. $\lor$ and $\land$ represent logical OR and AND, respectively. The symbol $\cdots$ indicates additional patterns that are omitted for brevity.
We ground the discriminative rules $\hat{\phi}_{M}$ and present the selective biochemical knowledge extracted by LogicXGNN, as well as by the baseline approaches GLGExplainer and GraphTrail, for the Mutagenicity and BBBP datasets in Figure 2. Additional results are provided in Appendix A.10.
First, it is worth noting that GLGExplainer cannot independently extract knowledge. It relies on prior domain knowledge to learn concept representations for each cluster of local explanations, which makes it less practical for uncovering unknown knowledge relevant to GNNs’ predictions. Furthermore, we notice both baselines require one-hot encoded inputs. In contrast, our grounding approach makes no assumptions about the input format and even supports continuous features.
Second, we observe that the substructures extracted by LogicXGNN are chemically accurate and scientifically meaningful, effectively explaining the classification outcomes. For example, we find that molecules containing oxygen-rich functional groups are less likely to cross the blood-brain barrier (Class 0 in BBBP), which can be attributed to their increased hydrophilicity, reduced lipophilicity, and greater likelihood of recognition by efflux transporters [17]. In the context of mutagenicity prediction (with class 0 indicating mutagenic compounds), we not only recover the well-known nitro group (NO2) attached to aromatic rings—widely recognized for its association with DNA damage and mutagenesis [9, 6]—but also identify other substructures, such as the trichloromethyl group (–CCl3), which is considered a structural alert for mutagenicity in cheminformatics and toxicology [33].
In contrast, baseline methods tend to extract less relevant or scientifically unsubstantiated substructures, or they identify only a limited subset of patterns that fail to match the breadth and chemical validity of those discovered by LogicXGNN. For example, GLGExplainer generates merely an $N\!-\!N\!-\!N$ substructure for Class 0 in Mutagenicity, accounting for just 38 of 2,463 instances. Similarly, GraphTrail generates just one pattern for Class 1. This limitation likely explains the observed performance gap between the baseline approaches and LogicXGNN, as shown in Table 1.
Figure 3: Selected examples generated by LogicXGNN and explanation graphs generated by XGNN. Each cell under LogicXGNN presents the original graph, its corresponding descriptive rules, and a newly generated instance created by modifying those rules. More details are in Appendix A.11.
@@ 4.3 How effective is $\phi_{M}$ as a generative model?
As the descriptive rules make each decision step in GNNs transparent, they enable accurate modeling of the underlying data distribution and fully controlled instance generation. Specifically, we can select a set of predicates, construct a graph with relevant connectivity patterns learned from the dataset, and apply grounding rules to assign node features, resulting in the final graph. More details on the graph generation process are provided in Appendix A.11. To demonstrate the above merits, we present examples generated by LogicXGNN and compare the explanation graphs generated by the baseline approach, XGNN, in Figure 3. Figure 9 (Appendix A.11) highlights the diversity in graph generation.
It is worth mentioning that our approach generates new instances that preserve similar structures and key properties, while ensuring adherence to chemical principles such as bonding accuracy. In contrast, XGNN employs a reinforcement learning agent that is inherently black-boxed, producing graphs that do not align with the actual data distribution. For instance, it generates bipartite structures that lack molecular relevance. In summary, LogicXGNN shows significant potential in fields such as drug design, and we plan to explore these possibilities further in future work.